# Shuofei Qiao♠♡, Pengjun Xie, Fei Huang, Huajun Chen, Ningyu Zhang

### v2

@@ Abstract
Large Language Models (LLMs) based agents excel at diverse tasks, yet they suffer from brittle procedural memory that is manually engineered or entangled in static parameters. In this work, we investigate strategies to endow agents with a learnable, updatable, and lifelong procedural memory. We propose $Mem^{p}$ that distills past agent trajectories into both fine-grained, step-by-step instructions and higher-level, script-like abstractions, and explore the impact of different strategies for Build, Retrieval, and Update of procedural memory. Coupled with a dynamic regimen that continuously updates, corrects, and deprecates its contents, this repository evolves in lockstep with new experience. Empirical evaluation on TravelPlanner and ALFWorld shows that as the memory repository is refined, agents achieve steadily higher success rates and greater efficiency on analogous tasks. Moreover, procedural memory built from a stronger model retains its value: migrating the procedural memory to a weaker model yields substantial performance gains.
Introduction
As large language models (LLMs) grow ever more powerful, LLM-based agents augmented by their own reasoning and external tools are taking on increasingly sophisticated works (Zhao et al. 2023; Wang et al. 2024a; Xi et al. 2025; Qiao et al. 2023). No longer mere assistants, these agents now trawl the web for elusive insights and weave them into comprehensive, publication ready reports, like Deep Research (OpenAI 2025; x.ai 2025) and WebDancer (Wu et al. 2025a). Moreover, they can handle complex data analyses (Lan et al. 2025; Ifargan et al. 2025; Ou et al. 2025), navigate multi-step GUI workflows (Luo et al. 2025; Qin et al. 2025), and sustain long-horizon, tool-rich interactions (Yao et al. 2025; Barres et al. 2025; Chen et al. 2025; Fang et al. 2025; Gur et al. 2023) with precision. Yet executing such intricate, long-horizon tasks demands dozens of steps and protracted runtimes. Along the way, unpredictable external events—network glitches, UI changes, shifting data schemas—can derail the entire process. Restarting from scratch every time is a punishing ordeal for present-day agents. Beneath their surface diversity, many complex tasks share deep structural commonalities and a similar environment. Instead of starting fresh each time, an agent should extract its experience from past successes. By turning earlier trajectories into reusable templates like patterns of reasoning, tool sequences, and recovery tactics, it can progress step by step, learning from every failure and success, until even the most convoluted missions become routine.
@@ Introduction
The capacity to distill, chronicle, and re-apply lessons from one’s own experiential trajectory is the bedrock of human learning and the pivotal gateway through which an agent ascends toward self-directed refinement (Liu et al. 2025a; Sumers et al. 2023a; Li et al. 2023). Procedural memory (Gupta and Cohen 2002; Cohen and Squire 1980) silently compiles habitual skills into executable subroutines, enabling unconscious, fluent action. While contemporary agents built on LLMs can compose short action plans or call external tools, their procedural knowledge is either hand-crafted, stored as brittle prompt templates, or implicitly entangled in model parameters that are expensive to update. Existing memory-augmented frameworks such as LangGraph (Mavroudis 2024), AutoGPT (Yang, Yue, and He 2023), or agent cognitive architectures like Memory Bank (Zhong et al. 2024a; Sumers et al. 2023b) and Soar (Laird 2022) provide coarse abstractions (buffers, rule chunks, production systems) but leave the optimization of procedural memory life-cycle operations about how skills are built, indexed, patched, and eventually pruned, largely unexamined. Consequently, there is no principled way to quantify how efficiently an agent evolves its procedural repertoire or to guarantee that new experiences improve rather than erode performance.
@@
We instantiate $Mem^{p}$ on top of strong LLMs (GPT-4o and Claude, Qwen) and evaluate on two diverse domains: long-horizon housework ALFWorld (Shridhar et al. 2021) and long-term information seeking task TravelPlanner (Xie et al. 2024). On two benchmark datasets that rigorously evaluate agent capabilities, we demonstrate that constructing and retrieving procedural memory during training empowers an agent to distill and reuse its prior experience. When this memory is exploited at test time, the agent’s task accuracy rises, and compared with tackling each instance in isolation, it eliminates most fruitless exploration on unfamiliar tasks, yielding substantial reductions in both step count and token consumption. Further, by equipping the agent with a set of memory-update mechanisms, we allow it to build and refine its procedural memory while acting in the test environment. This endows the agent with a continual, almost linear mastery of the task. Extensive ablations reveal that procedural memory also scales gracefully and transfers effectively to new, related tasks.
Related Works
@@
Memory is a foundational component in language agents, enabling them to retain and utilize past information across multiple timescales, including short-term, episodic, and long-term memory, to enhance their performance and adaptability (Zhou et al. 2023, 2024; Zhang et al. 2024; Liu et al. 2025a; Li et al. 2025). These systems aim to mimic aspects of human memory to improve coherence, personalization, and learning capabilities  (Chhikara et al. 2025; Wu et al. 2025b; Xia et al. 2025). Current approaches include end-to-end memory systems (Yu et al. 2025; Zhou et al. 2025), external memory systems (Chhikara et al. 2025; Zhong et al. 2024b), and hierarchical memory structures (Hu et al. 2024a; Xu et al. 2025). These methods involve encoding and storing information in various formats, using retrieval mechanisms like vector embeddings and semantic search, and implementing memory updating and forgetting strategies to maintain relevance and efficiency. Despite its importance, memory in multi-turn agent interactions remains underexplored, and enabling agents to effectively learn and utilize memory across trajectories poses a significant challenge. Procedural memory is a type of long-term memory that involves the retention of procedures and skills, such as typing or riding a bike, which are performed automatically without conscious thought. The agent utilizes procedural memory to internalize and automate repetitive tasks, decision-making processes, and interaction patterns, leading to more efficient and context-aware responses over time. Although there have been several works, such as Voyager (Wang et al. 2023), AWM (Wang et al. 2024b), and AutoManual (Chen et al. 2024), that utilize procedural memory to enhance agents’ capabilities on similar tasks, there still lacks a systematic analysis on how to construct, retrieve, and update such procedural memory like  (Wu et al. 2024). Therefore, our work mainly focuses on exploring how to build an effective procedural memory system for agents performing cross-trajectory tasks.
@@
LLM-based Agent learning from experience involves intelligence continuously improving their decision-making capabilities through interaction with environments and utilization of past experiences (Tan et al. 2025; Tang et al. 2025; Zhou et al. 2025; Qiao et al. 2025; Su et al. 2025; Wang et al. 2024b). This approach is crucial for developing adaptive and intelligent agents capable of handling dynamic real-world scenarios, as it allows them to optimize behaviors, reduce manual programming needs, and enhance performance across various tasks (Zheng et al. ; Liu et al. 2025c; Wang et al. 2025). Agents typically employ mechanisms such as reinforcement learning (Lu et al. 2025; Dong et al. 2025), experience replay (Feng et al. 2025; Liu et al. 2025b), imitation learning (Sun et al. 2024; Yang et al. 2024b), memory management (Hou, Tamoto, and Miyashita 2024; Hu et al. 2024b), and multi-agent learning to achieve this. However, current methods face limitations including low sample efficiency, poor generalization across tasks, catastrophic forgetting when learning new information, and there are very few features for memory update. The key distinction of our work lies in systematically investigating optimal strategies for construction, retrieval, and update modules of an agent’s procedural knowledge. During the update phase, we enhance the agent’s capabilities by maintaining an editable repository of procedural knowledge. Additionally, collecting high-quality training data can be challenging and may introduce biases. Addressing these limitations is essential for advancing the capabilities of LLM-based agents and ensuring their effective application in real-world contexts.
Preliminary
@@
When an agent influences its external environment by invoking external tools or executing prescribed actions, and iteratively refines its behavior over multiple rounds to accomplish a complex multi-step objective, this paradigm can be modeled as a Markov Decision Process (MDP). (Puterman 1990) Under this view, at each discrete time step $t$, the agent, situated in state $s_{t}\in S$, chooses an action $a_{t}\in A$, according to its policy $\pi(a_{t}|s_{t})$, where $A$ is the action space of the task. The environment then transitions to a new state $s_{t+1}\in S$ and emits an observation $O_{t}$. Consequently, the entire interaction trajectory may be compactly expressed as:
$\tau=(s_{0},a_{0},o_{1},s_{1},a_{1},o_{2},\ldots,s_{T}),$ (1)
where $\tau$ is the complete exploration trajectory of this task. Moreover, a reward function $R$ will evaluate the task’s completion $r$ within this environment envenv by assigning a score based on the final state $s_{T}$ or the entire trajectory $\tau$.
$r=R(env,s_{T},\tau)\in[0,1]$ (2)
@@
By shifting from parallel task completion to sequential task completion, the agent can learn and distill experience from earlier tasks, thereby reducing repetitive exploration. Inspired by human procedural memory, we propose to equip the agent with a procedural memory module. This module transforms the conventional policy $\pi(a_{t}|s_{t})$ into $\pi_{m^{p}}(a_{t}|s_{t})$, where $m^{p}$ is the agent’s learned procedural memory.
Agent Procedural Memory
Procedural memory is the type of long-term memory responsible for knowing how to perform tasks and skills, such as typing or riding a bike. By mastering this type of procedural memory, humans avoid the need to relearn the process each time. For an agent, that is, for a task trajectory $\tau$ and its reward $r$, a memory $m^{p}$ is constructed by a builder $B$, thereby achieving the acquisition of memory, namely
$Mem=\sum_{t=1}^{T}m^{p_{t}},where\ m^{p_{t}}=B(\tau_{t},r_{t})$ (3)
@@
As we use cosine similarity for the vector embedding model $\phi$ of the task in the experiment, the retrieval process becomes:
@@
Let $M(t)$ denote the agent’s procedural memory at time $t$, and $\tau_{t}$ represent the set of tasks completed up to time $t$. Then, the update mechanism can be modeled as a function $U$ that takes the current procedural memory and task execution feedback to produce the updated memory:
$M(t+1)=U\left(M(t),E(t),\tau_{t}\right),$ (6)
@@
$U=Add\left(M_{new}\right)\ominus Remove\left(M_{obso}\right)\oplus Update(M_{exist}),$ (7)
where $M_{new}$ represents new procedural memory to be added; $M_{obso}$ indicates procedural memory to be removed, $M_{existing}$ are tasks to be updated based on execution feedback $E(t)$. This comprehensive formula captures the essential add, delete, and modify operations within the update mechanism.
Experiment
@@
#CS #HC Steps ↓ Dev ↑ Test ↑ Steps ↓
@@
Table 1: Results on Build Policy. #CS, #HC denote Commensense and Hard Constraint, respectively. ↑ indicates the higher values are better, and ↓ denotes the lower values are better. The best results among all methods with similar settings are bolded, and the second-best results are underlined.
In this section, we will introduce the Procedural Memory framework in detail (Figure 2), covering the storage, retrieval, and update modules of memory, as well as analyzing which strategies perform better within each module.
Experimental Settings
@@
For our experiments, we adapt TravelPlanner (Xie et al. 2024) and ALFWorld (Shridhar et al. 2021) benchmarks. TravelPlanner is a benchmark designed to evaluate agents’ ability to use tools and perform complex planning under intricate constraints. In contrast, ALFWorld comprises household tasks. In each interaction round, the agent outputs an action, and the environment responds with textual feedback describing the resulting state. This process repeats for multiple turns until the task is completed or the maximum number of rounds is reached. ALFWorld includes test split to evaluate the agent’s generalization ability.
@@
In our experiments, we benchmarked our procedural memory on three base models. Specifically, we adopt the two proprietary frontier models that have consistently dominated public leaderboards: OpenAI’s GPT-4o (OpenAI 2022) and Anthropic’s Claude (Anthropic 2022), and complement them with the open-sourced Qwen2.5-72B-Instruct (Yang et al. 2024a). The first two provide state-of-the-art closed-source performance, while the third allows us to verify that our findings generalize beyond proprietary systems and remain valid in the open-source regime.
Evaluation.
Memory Storage & Retrieval
@@
As shown in Table 1, all memory construction methods outperform the no-memory baseline, achieving higher scores on both datasets while also reducing the number of steps required. This indicates that procedural memory built during training is beneficial for directly applying tasks during testing. Furthermore, we observe that the approach of abstracting trajectories into scripts during training yields relatively better performance on the ALFWorld test set compared to the dev set. Conversely, trajectories that utilize complete execution traces as procedural memory achieve higher scores on the dev set, suggesting that scripts are more capable of generalizing to different test tasks, while trajectories are better suited for scenarios involving tasks similar to those already completed. By combining procedure knowledge from both methods of employing abstracted guidelines along with concrete execution trajectories, we attain the optimal performance.
@@
During the retrieval process, we evaluate the similarity by calculating the cosine similarity between their corresponding vectors. Our experiments show that these different retrieval strategies produce varying results. Specifically, compared to random sampling, employing the query based and AveFact methods for precise retrieval significantly improves performance. The query-based approach benefits from capturing semantic contexts, enabling more accurate matches. The AveFact method, by extracting key features and averaging their similarities, effectively focuses on core task elements, leading to better retrieval efficacy. Overall, our findings suggest that incorporating semantic understanding and key feature extraction in retrieval strategies substantially enhances memory access accuracy and the effectiveness of downstream task performance.
Model Policy #CS #HC Steps ↓
@@
Memory Update
While many prior efforts have focused on developing reusable procedural knowledge, enabling models to learn from prior experiences rather than solving each test task in isolation, most existing memory update methods remain quite rudimentary. Typically, they simply append newly acquired memories to the existing store—a so-called ”merge” strategy. In this work, we explore several online memory-update mechanisms to identify which dynamic strategy delivers the best performance on our tasks. Beyond end-to-end evaluation metrics, we also analyze how both accuracy and efficiency evolve as the number of executed tasks increases, explicitly measuring the benefits conferred by our procedural memory.
To facilitate systematic comparison, we designed several memory-update scenarios. In each, the agent’s episodic memory is refreshed after every t test-set tasks. The specific update strategies are as follows:

### v3

@@ Abstract
Large Language Models (LLMs) based agents excel at diverse tasks, yet they suffer from brittle procedural memory that is manually engineered or entangled in static parameters. In this work, we investigate strategies to endow agents with a learnable, updatable, and lifelong procedural memory. We propose $Mem^{p}$ that distills past agent trajectories into both fine-grained, step-by-step instructions and higher-level, script-like abstractions, and explore the impact of different strategies for Build, Retrieval, and Update of procedural memory. Coupled with a dynamic regimen that continuously updates, corrects, and deprecates its contents, this repository evolves in lockstep with new experience. Empirical evaluation on TravelPlanner and ALFWorld shows that as the memory repository is refined, agents achieve steadily higher success rates and greater efficiency on analogous tasks. Moreover, procedural memory built from a stronger model retains its value: migrating the procedural memory to a weaker model can also yield substantial performance gains.
$Mem^{p}$: Exploring Agent Procedural Memory
[…]
1 Introduction
As large language models (LLMs) grow ever more powerful, LLM-based agents augmented by their own reasoning and external tools are taking on increasingly sophisticated works Zhao etĀ al. (2023); Wang etĀ al. (2024a); Xi etĀ al. (2025); Qiao etĀ al. (2023). No longer mere assistants, these agents now trawl the web for elusive insights and weave them into comprehensive, publication ready reports, like Deep ResearchĀ OpenAI (2025); x.ai (2025) and WebDancerĀ Wu etĀ al. (2025a). Moreover, they can handle complex data analysesĀ Lan etĀ al. (2025); Ifargan etĀ al. (2025); Ou etĀ al. (2025), navigate multi-step GUI workflowsĀ Luo etĀ al. (2025); Qin etĀ al. (2025), and sustain long-horizon, tool-rich interactionsĀ Yao etĀ al. (2025); Barres etĀ al. (2025); Chen etĀ al. (2025); Fang etĀ al. (2025); Gur etĀ al. (2023) with precision. Yet executing such intricate, long-horizon tasks demands dozens of steps and protracted runtimes. Along the way, unpredictable external events—network glitches, UI changes, shifting data schemas—can derail the entire process. Restarting from scratch every time is a punishing ordeal for present-day agents. Beneath their surface diversity, many complex tasks share deep structural commonalities and a similar environment. Instead of starting fresh each time, an agent should extract its experience from past successes. By turning earlier trajectories into reusable templates like patterns of reasoning, tool sequences, and recovery tactics, it can progress step by step, learning from every failure and success, until even the most convoluted missions become routine.
@@ 1 Introduction
The capacity to distill, chronicle, and re-apply lessons from one’s own experiential trajectory is the bedrock of human learning and the pivotal gateway through which an agent ascends toward self-directed refinementĀ Liu etĀ al. (2025a); Sumers etĀ al. (2023a); Li etĀ al. (2023). Procedural memoryĀ Gupta and Cohen (2002); Cohen and Squire (1980) silently compiles habitual skills into executable subroutines, enabling unconscious, fluent action. While contemporary agents built on LLMs can compose short action plans or call external tools, their procedural knowledge is either hand-crafted, stored as brittle prompt templates, or implicitly entangled in model parameters that are expensive to update. Existing memory-augmented frameworks such as LangGraphĀ Mavroudis (2024), AutoGPTĀ Yang etĀ al. (2023), or agent cognitive architectures like Memory BankĀ Zhong etĀ al. (2024a); Sumers etĀ al. (2023b) and SoarĀ Laird (2022) provide coarse abstractions (buffers, rule chunks, production systems) but leave the optimization of procedural memory life-cycle operations about how skills are built, indexed, patched, and eventually pruned, largely unexamined. Consequently, there is no reliable way to measure how efficiently an agent grows its procedural skills or to ensure new experiences improve rather than harm performance.
@@
We instantiate $Mem^{p}$ on top of strong LLMs (GPT-4o and Claude, Qwen) and evaluate on two diverse domains: long-horizon housework ALFWorldĀ Shridhar etĀ al. (2021) and long-term information seeking task TravelPlannerĀ Xie etĀ al. (2024). On two benchmark datasets that rigorously evaluate agent capabilities, we demonstrate that constructing and retrieving procedural memory during training empowers an agent to distill and reuse its prior experience. When this memory is exploited at test time, the agent’s task accuracy rises, and compared with tackling each instance in isolation, it eliminates most fruitless exploration on unfamiliar tasks, yielding substantial reductions in both step count and token consumption. Further, by equipping the agent with a set of memory-update mechanisms, we allow it to build and refine its procedural memory while acting in the test environment. This endows the agent with a continual, almost linear mastery of the task. Extensive ablation studies reveal that procedural memory not only scales gracefully with increasing task complexity but also transfers effectively to new, related tasks, demonstrating its adaptability across scenarios.
2 Related Works
@@
Memory is a foundational component in language agents, enabling them to retain and utilize past information across multiple timescales, including short-term, episodic, and long-term memory, to enhance their performance and adaptabilityĀ Zhou etĀ al. (2023, 2024); Zhang etĀ al. (2024); Liu etĀ al. (2025a); Li etĀ al. (2025). These systems aim to mimic aspects of human memory to improve coherence, personalization, and learning capabilities Ā Chhikara etĀ al. (2025); Wu etĀ al. (2025b); Xia etĀ al. (2025). Current approaches include end-to-end memory systemsĀ Yu etĀ al. (2025); Zhou etĀ al. (2025), external memory systemsĀ Chhikara etĀ al. (2025); Zhong etĀ al. (2024b), and hierarchical memory structuresĀ Hu etĀ al. (2024a); Xu etĀ al. (2025). These methods involve encoding and storing information in various formats, using retrieval mechanisms like vector embeddings and semantic search, and implementing memory updating and forgetting strategies to maintain relevance and efficiency. Despite its importance, memory in multi-turn agent interactions remains underexplored, and enabling agents to effectively learn and utilize memory across trajectories poses a significant challenge. Procedural memory is a type of long-term memory that involves the retention of procedures and skills, such as typing or riding a bike, which are performed automatically without conscious thought. The agent utilizes procedural memory to internalize and automate repetitive tasks, decision-making processes, and interaction patterns, leading to more efficient and context-aware responses over time. Although there have been several works, such as VoyagerĀ Wang etĀ al. (2023), AWMĀ Wang etĀ al. (2024b), and AutoManualĀ Chen etĀ al. (2024), that utilize procedural memory to enhance agents’ capabilities on similar tasks, there still lacks a systematic analysis on how to construct, retrieve, and update such procedural memory like Ā Wu etĀ al. (2024). Therefore, our work mainly focuses on exploring how to build an effective procedural memory system for agents performing cross-trajectory tasks.
@@
LLM-based Agent learning from experience involves intelligence continuously improving their decision-making capabilities through interaction with environments and utilization of past experiencesĀ Tan etĀ al. (2025); Tang etĀ al. (2025); Zhou etĀ al. (2025); Qiao etĀ al. (2025); Su etĀ al. (2025); Wang etĀ al. (2024b). This approach is crucial for developing adaptive and intelligent agents capable of handling dynamic real-world scenarios, as it allows them to optimize behaviors, reduce manual programming needs, and enhance performance across various tasksĀ Zheng etĀ al. ; Liu etĀ al. (2025c); Wang etĀ al. (2025). Agents typically employ mechanisms such as reinforcement learningĀ Lu etĀ al. (2025); Dong etĀ al. (2025), experience replayĀ Feng etĀ al. (2025); Liu etĀ al. (2025b), imitation learningĀ Sun etĀ al. (2024); Yang etĀ al. (2024b), memory managementĀ Hou etĀ al. (2024); Hu etĀ al. (2024b), and multi-agent learning to achieve this. However, current methods face limitations including low sample efficiency, poor generalization across tasks, catastrophic forgetting when learning new information, and there are few features for memory update. The key distinction of our work lies in systematically investigating optimal strategies for construction, retrieval, and update modules of an agent’s procedural knowledge. During the update phase, we enhance the agent’s capabilities by maintaining an editable repository of procedural knowledge. Additionally, collecting high-quality training data can be challenging and may introduce biases. Addressing these limitations is essential for advancing the capabilities of LLM-based agents and ensuring their effective application in real-world contexts.
3 Preliminary
@@
When an agent influences its external environment by invoking external tools or executing prescribed actions, and iteratively refines its behavior over multiple rounds to accomplish a complex multi-step objective, this paradigm can be modeled as a Markov Decision Process (MDP).Ā Puterman (1990) Under this view, at each discrete time step $t$, the agent, situated in state $s_{t}\in S$, chooses an action $a_{t}\in A$, according to its policy $\pi(a_{t}|s_{t})$, where $A$ is the action space of the task. The environment then transitions to a new state $s_{t+1}\in S$ and emits an observation $O_{t}$. Consequently, the entire interaction trajectory may be compactly expressed as:
[…]
@@
[…]
3.1 Agent Procedural Memory
[…]
@@ 3.1 Agent Procedural Memory
[…]
@@
[…]
@@
$U=Add\left(M_{new}\right)\ominus Del\left(M_{obs}\right)\oplus Update(M_{est})$ (7)
where $M_{new}$ represents new procedural memory to be added; $M_{obs}$ indicates procedural memory to be removed, $M_{est}$ are tasks to be updated based on execution feedback $E(t)$. This comprehensive formula captures the essential add, delete, and modify operations within the update mechanism.
4 Experiment
@@
#CS ↑ #HC ↑ Steps ↓ Dev ↑ Test ↑ Steps ↓
@@
Table 1: Results on Build Policy. #CS, #HC denote Commonsense and Hard Constraint, respectively. ↑ indicates the higher values are better, and ↓ denotes the lower values are better. The best results among all methods with similar settings are bolded, and the second-best results are underlined.
In this section, we will introduce the Procedural Memory framework in detail (FigureĀ 2), covering the storage, retrieval, and update modules of memory, as well as analyzing which strategies perform better within each module.
4.1 Experimental Settings
@@
For our experiments, we adapt TravelPlannerĀ (Xie etĀ al., 2024) and ALFWorldĀ (Shridhar etĀ al., 2021) benchmarks. TravelPlanner is a benchmark designed to evaluate agents’ ability to use tools and perform complex planning under intricate constraints. In contrast, ALFWorld comprises household tasks. In each interaction round, the agent outputs an action, and the environment responds with textual feedback describing the resulting state. This process repeats for multiple turns until the task is completed or the maximum number of rounds is reached. ALFWorld includes test split to evaluate the agent’s generalization ability. Detailed evaluation methods are shown in AppendixĀ C.
@@
In our experiments, we benchmarked our procedural memory on three base models. Specifically, we adopt the two proprietary frontier models that have consistently dominated public leaderboards: OpenAI’s GPT-4oĀ OpenAI (2022) and Anthropic’s ClaudeĀ Anthropic (2022), and complement them with the open-sourced Qwen2.5-72B-InstructĀ Yang etĀ al. (2024a). The first two provide state-of-the-art closed-source performance, while the third allows us to verify that our findings generalize beyond proprietary systems and remain valid in the open-source regime.
4.2 Memory Storage & Retrieval
@@
As shown in TableĀ 1, all memory construction methods outperform the no-memory baseline, achieving higher scores on both datasets while also reducing the number of steps required. This indicates that procedural memory built during training is beneficial for directly applying tasks during testing. Furthermore, we observe that the approach of abstracting trajectories into scripts during training yields relatively better performance on the ALFWorld test set compared to the dev set. Conversely, trajectories that utilize complete execution traces as procedural memory achieve higher scores on the dev set, suggesting that scripts are more capable of generalizing to different test tasks, while trajectories are better suited for scenarios involving tasks similar to those already completed. By combining procedure knowledge from both methods of employing abstracted guidelines along with concrete execution trajectories, we attain the optimal performance.
@@
During the retrieval process, we evaluate the similarity by calculating the cosine similarity between their corresponding vectors as shown in Table 2. Our experiments show that these different retrieval strategies produce varying results. Specifically, compared to random sampling, employing the query based and AveFact methods for precise retrieval significantly improves performance. The query-based approach benefits from capturing semantic contexts, enabling more accurate matches. The AveFact method, by extracting key features and averaging their similarities, effectively focuses on core task elements, leading to better retrieval efficacy. Overall, our findings suggest that incorporating semantic understanding and key feature extraction in retrieval strategies substantially enhances memory access accuracy and the effectiveness of downstream task performance.
Model Policy #CS ↑ #HC ↑ Steps ↓
@@
4.3 Memory Update
While many prior efforts have focused on developing reusable procedural knowledge, enabling models to learn from prior experiences rather than solving each test task in isolation, most existing memory update methods remain quite rudimentary. Typically, they simply append newly acquired memories to the existing store—a so-called ā€œmergeā€ strategy. In this work, we explore several online memory-update mechanisms to identify which dynamic strategy delivers the best performance on our tasks. Beyond end-to-end evaluation metrics, we also analyze how both accuracy and efficiency evolve as the number of executed tasks increases, explicitly measuring the benefits conferred by our procedural memory.
To facilitate systematic comparison, we design several memory-update scenarios. The agent’s episodic memory is refreshed after every t test-set tasks, following the specific update strategies: